OpenStudy (anonymous):

when finding the positive acute angle of an angle that lies in quadrant 3, do you subtract thetha from 180?

OpenStudy (anonymous):

*180 from thetha

OpenStudy (anonymous):

Em I am not really sure about this one can you help

OpenStudy (anonymous):

can i give you the problem

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

cos 230 as a positive acute angle?

OpenStudy (anonymous):

i think its actually - cos50 but idk how to do the work

OpenStudy (anonymous):

I dont remember well but there was a principle cos(-a) = cos(a)

jigglypuff314 (jigglypuff314):

erm ... the angle 230 makes with the x axis would be 50 degrees because 230-180 = 50 then the "All Students Take Calculus" thing https://en.wikipedia.org/wiki/All_Students_Take_Calculus so you know cos is negative in Quadrant 3

jigglypuff314 (jigglypuff314):

so yeah naruto is right :)

OpenStudy (anonymous):

omg Em you are intelligent ^_^

jigglypuff314 (jigglypuff314):

tan(-45) = tan(360-45) = tan(315) and 315 is just 45 under the x axis (quadrant 4) tan is negative in Q4 so so -tan(45) .... which can be figured out with a 45 45 90 triangle

OpenStudy (anonymous):

wait.... i don't get it.....

OpenStudy (anonymous):

i cant even figure out the answer.. isnt it supposed to be acute?

jigglypuff314 (jigglypuff314):

sorry >.< annas asked me to help, but I'm just a bit tired XD so tan(45) = 1 tan(-45) is in quadrant 4 so tan(-45) = -1

OpenStudy (anonymous):

i realize that so wont the answer be similar like - tan 45... can that even be right?/

jigglypuff314 (jigglypuff314):

yeah tan(-45) = -tan(45) because tan(-45) = tan(315) = -tan(45)

OpenStudy (anonymous):

but can it be the same? is that right... i mean its basically the same thing..

jigglypuff314 (jigglypuff314):

mmm yeah :) but but I'm not sure if it always works like that, so be careful

OpenStudy (anonymous):

thanks so much ! i hate take home tests -_-

jigglypuff314 (jigglypuff314):

LOL I always seem to mess them up some how XD